Minutes of Working Group Meeting


Subaru Base Facility, Hilo

March 13, 2014

Present:

Absent:

Minutes

  1. Current Network Status
  2. On January 18, the SX south path was down between 9:30 and 12:00 UTC for the maintenance of one of its core nodes in Los Angeles Big Island Internet2 traffic was rerouted through Oahu. No issues were reported.

    On Sunday February 1, between 7:30am and 8:00am, UH-ITS took down the entire UH network on Oahu to install the Palo Alto network device, a next generation firewall that could identify and provide fine-grained control of applications and an intrusion prevention system that identifies applications regardless of port, protocol, evasive tactic or Secure Sockets Layer at 10 Gbps throughput.

    Soon after the installation various problems were observed by the Maunakea observatories and UH Hilo. The problems were mostly related to web applications. UH-ITS was contacted and they traced the cause of problems to the mishandling of multiple links and asymetric paths by the Palo Alto device. Adjustments were made and all known issues were fixed by February 4.

    On February 5, observatories using the IfA Domain Name Server (DNS) were observing slow responses. The problem was intermittent. Suspecting a DOS attack on the IfA DNS, Pui Hin asked the affected observatories and users to temporarily switch to the UH-ITS DNS. Issues with the IfA DNS were later solved by further restricting recursive lookups to known sites only.

  3. Upgrade of the Observatories' Connections to the MKOCN
  4. Pui Hin reported that the alternate 1 Gbps observatories Internet connection in Hilo and on the Summit have been configured. As a reminder, with the primary route now supported by a Cisco 4500x, the old Cisco 3750Gs were reconfigured to provide the observatories with a backup Internet/Internet2 connection. Observatories are invited to contact Pui Hin or Miranda to upgrade their primary connections to 10 Gbps and to set up their backup links.

  5. Weather Display Station at HP
  6. Pui Hin has confirmed with Stewart Hunter that there is money within the current year's budget to replace the HP weather station and will go ahead with the purchase of the computer, monitors, and mounting brackets.

    Meanwhile Miranda has been working with the observatories on their contributions which include codes, Perl scripts, system designs, and system scripts. She also has had some success in adopting the codes and scripts contributed. However her progress has been restricted by the limited old equipment that she was using. She is expected to be able to make further progress once the new equipment becomes available. During the rest of the meeting, Miranda worked with the group to finalize the preliminary set of displays.

  7. Next meeting
  8. June 12th, Thursday, 10:00am at the IfA Hilo Facility.


MKOCN Homepage Back to Minutes

Last updated March 20, 2013 by Miranda Hawarden-Ogata